# Automated Black-box Prompt Engineering for Personalized Text-to-Image Generation

### v1

@@ 3.4 Designing the judge model $\mathbf{D}$
We have a wider range of choices for the judging model as long as it provides a notion of similarity between a pair of images. A simple solution is to use pre-trained discriminative models such as CLIP [22] and DINO [21], and measure the distance of images in their embedding spaces. These models have seen various degrees of success but come with inherent limitations – the discriminative objective (e.g., contrastive loss) does not incentivize the model to attend to fine-grained details since they do not improve the objective further, an issue similar to the shortcomings of using captioning models to generate prompts [13]. Moreover, in image generation, the criteria of success can be nuanced and difficult to quantify through traditional distance or similarity functions yet can be effortlessly described in human language. Lastly, the similarity we wish to measure may only involve some part of the visual features (e.g. color, painting stroke type, etc), and not all applications share the same notion of similarity. If we want to use pre-trained discriminative models, then we need to find a different model for each specific task, which can be impractical.
In light of these challenges, an ideal judge model should be maximally flexible for different kinds of criteria and can perform fine-grained analysis of the images. Once again, a multimodal LLM emerges as the perfect candidate: using system prompts and in-context learning, we can easily specify metrics that may be otherwise difficult to describe or evaluate and even intervene in the reasoning chain if we want to, and, more importantly, the same model can be applied to a wide range of tasks.
@@ 4.1 Experimental Settings
4.1.1 Implementation Details
For all of our experiments, we choose GPT-4V [20] as both the prompt engineer assistant model ${\mathbf{F}}$ and the judge ${\mathbf{D}}$. We also fix the T2I generator as SDXL-Turbo [27] for all of our experiments. We design different system prompts for both ${\mathbf{F}}$ and ${\mathbf{D}}$ for each task and we provide details about the system prompts in the appendix.
@@ 4.1.1 Implementation Details
We compare PRISM and baselines in two settings: personalized T2I generation and direct image inversion, and we will elaborate on the task definitions in their corresponding sections below. For personalized T2I generation, we use a maximum budget of $40$ and report the quantitative results from the setting $N=10,K=4$. For direct image inversion, we use a maximum budget of $30$ and report the quantitative results from the setting $N=6,K=5$. To simplify the implementation, we only keep a chat history length of $3$ and use the length of the prompt as an approximation of the prompt log-likelihood in the final prompt selection. For direct image inversion, we re-evaluate the top $5$ candidates twice and tally the score with the in-iteration scores to make the final decision. For personalized T2I generation, we re-evaluate once for each reference image and use the average score to select the output.
4.1.2 Baselines
@@
4.1.3 Evaluation Metrics
We evaluate the prompt interpretability using mean negative log-likelihood (NLL) calculated from Mistral 7B [11]. For image quality evaluation, we mainly measure the CLIP image similarity score (CLIP-I) to quantify the difference between the generated images and the reference images. Following [26], we also use DINO V2 [21] embedding similarity to calculate the object-sensitive image similarity for the personalized T2I generation task. We chose CLIP-ViT-L-14 and DINO-V2-Base as the base models. For Dall-E 2 and Dall-E 3, we also compare the number of times each method fails to pass its black-box safeguard. More failures indicate a higher potential to produce unsafe prompts. For each prompt, we allow $5$ attempts before counting it as a failure.
@@ 4.2 Personalized Text-to-Image Generation
We first demonstrate PRISM’s ability to find human-interpretable and accurate prompts to describe certain objects and styles in the task of personalized T2I generation. Given a set of reference images that depict the same concept (such as objects and style), personalized T2I tasks require the model to synthesize images in new contexts while maintaining the original concept.
4.2.1 Datasets
We use the dataset collected by DreamBooth [26] to quantitatively compare the performance in personalized T2I generation. The DreamBooth dataset contains 30 daily objects, and each subject has 4-6 images. For each subject, we adopt the 25 prompt templates curated by DreamBooth to create varying contexts and scenarios to test the fidelity of the subject representation in diverse settings. We generate 4 images for each subject and template combination with open sourced T2I models, and 1 image for each combination with closed sourced T2I models. For methods that directly find English prompts, we use the class noun to fill in the template and the output prompts that describe these concepts serve as suffixes. For Textual Inversion, we follow the original setting for the templates.
We also qualitatively demonstrate the ability to represent a certain artistic style using Wikiart dataset [34]. We use three images from each artist as reference images. To create diverse scenes, we follow [8] and use descriptive prompts from PartiPrompts [42] as prefixes to the output prompts similar to the previous setting.
4.2.2 DreamBooth Dataset Results
Table 1 and Figure 3 respectively show the quantitative and qualitative results on the DreamBooth dataset. As we can observe, PRISM achieves the best performance across the board except for the image similarity metrics for SDXL-Turbo. PRISM is the only method in our experiments that can produce fully human-readable prompts for these subjects. In particular, we can observe that PEZ renders completely indecipherable texts, BLIP-2 only describes the general scene but fails to mention any visual details and textual inversion is entirely not interpretable since it produces soft embeddings. Since CLIP-Interrogator combines the results from BLIP-2 and a CLIP search, it improves the interpretability over PEZ-like gradient search-only method. However, it still falls short in terms of human readability in comparison to our method.
In terms of image quality and object fidelity, we also find PRISM to constantly achieve accurate depiction of the target subject while the baselines sometimes struggle to capture all the details. Fine-grained details such as the color of the dog and cat fur and the shape of the shoe sole are better described and reflected with our method. And out of the four training-free methods we experiment with, PRISM is the only one that can tackle complicated objects such as the red monster toy and the dog-shaped backpack as shown in Figure 3 when all the other methods fail to generate similar objects. Due to the nature of their methodologies, BLIP-2 and CLIP-Interrogator also capture the background and other irrelevant elements in the scene when describing the objects. However, unlike our method, where we can directly specify the tasks and the judging criteria in the system prompts of the LLMs, there is no simple way to automatically filter out those irrelevant elements in BLIP-2 and CLIP-Interrogator’s outputs. Even though Textual Inversion obtains marginally higher CLIP-I scores and DINO scores with SDXL-Turbo, notice that Textual Inversion requires a lot more modeling assumptions than our method, and the new embeddings it learns are not transferable – not even to SD 2.1, because SDXL models use two text encoders whereas SD 2.1 model use only one.
When transferring the output prompts to black-box T2I models, our method shows even larger advantages over the baselines. We also observe that our method produces the fewest unsafe prompts judged by Dall-E safeguards, while the baselines can fail to pass the safeguard up to almost $16.7\%$ of the time.
4.2.3 Wikiart Results
@@
In Figure 4, we also show a qualitative comparison between our method and the baselines on the Wikiart dataset. We find that our method is capable of precisely identifying the genres, eras, and sometimes even the names of the artists when describing the style of the reference artworks. On the other hand, we observe that the baselines fail to recognize these crucial keywords, even when they have access to a pre-collected bank of words that are supposed to provide accurate descriptions of the style. In addition, PRISM can also provide other fine-grained details such as the style of the pen strokes and color palettes in a human-interpretable way to better assist the generation of the target style.
@@ 4.3 Direct Image Inversion
To demonstrate the versatility of our method, we also compare PRISM with the baselines in the task of direct image inversion. In this task, the goal is to directly find the prompt that can exactly generate the input image. Here the number of reference images is $M=1$ and we aim to capture all aspects of the image, including the subjects, background, theme, style, and other details in the scene.
4.3.1 Datasets
We use images from the DiffusionDB dataset [35] for the direct image inversion task. This dataset includes a wide variety of image pairs generated by Stable Diffusion and we choose a random sample of 100 images from the large_random_10k split on Huggingface.
4.3.2 Results
As shown in Table 2, we can first obtain a similar observation in terms of human-interpretability to the previous experiment. For direct image inversion, we also immediately see a significant improvement in the readability of inverted prompts using PRISM. While expected for methods, such as PEZ, in which the process of image inversion has no language prior, we find that our method finds text that more closely aligns with a learned distribution of English language text (i.e. Mistral logits) than CLIP-Interrogator and BLIP2.
When comparing the image quality, we first note that because all images in DiffusionDB are generated by Stable Diffusion, which is exactly the model design space of CLIP-Interrogator and PEZ, it gives significant modeling assumption advantages to these baselines over our method when testing on Stable Diffusion models. This advantage enables relatively high performance for these baselines in Stable Diffusion models, but it does not transfer well into other closed-sourced models. In fact, we can even observe that CLIP-Interrogator generates the highest quality images with SD 2.1, which is the weakest model in this comparison and generates the lowest quality images with Dall-E 3, which is the strongest T2I model in this table. This phenomenon indicates that the design choices of CLIP-Interrogate and PEZ are heavily overfitted to achieve high performance on Stable Diffusion, but provide poor generalizability to other models. On the other hand, the prompts produced by our method generalize significantly better than the baselines and we achieve the best results on Dall-E 3.
@@ 4.4 Ablation Study
Figure 6: Comparison with GPT-4V in both personalized T2I generation and direct image inversion experiments.

### v2

@@ 3.4 Designing the judge model $\mathbf{D}$
We have a wider range of choices for the judge model as long as it provides a notion of similarity between a pair of images. A simple solution is to use pre-trained discriminative models such as CLIP (Radford et al., 2021), and measure the distance in their embedding spaces. While these models have seen various degrees of success, they also come with inherent limitations – the discriminative objective (e.g., contrastive loss) does not incentivize the model to attend to fine-grained details, an issue similar to the shortcomings of using captioning models to generate prompts (Liang et al., 2023). Moreover, in image generation, the criteria of success can be nuanced and difficult to quantify through traditional similarity metrics yet can be effortlessly described in human language. Lastly, the similarity we wish to measure may only involve some part of the visual features (e.g. color), and not all applications share the same notion of similarity. If we want to use pretrained discriminative models, then we need to find a different model for each task, which can be inefficient.
In light of these challenges, an ideal judge model should be flexible for different kinds of criteria and can perform fine-grained analysis of the images. Once again, a VLM emerges as the perfect candidate: using system prompts and in-context learning, we can easily specify metrics that may be otherwise difficult to describe or evaluate and even intervene in the reasoning chain if we want to, and, more importantly, the same model can be applied to a wide range of tasks.
@@ 4.1 Experimental Settings
Implementation Details
For all of our experiments, we choose GPT-4V (OpenAI, 2023) as both the prompt engineer assistant model ${\mathbf{F}}$ and the judge ${\mathbf{D}}$. We also fix the T2I generator as SDXL-Turbo (Sauer et al., 2023) for all of our experiments. We design different system prompts for both ${\mathbf{F}}$ and ${\mathbf{D}}$ for each task and we provide details about the system prompts in the appendix.
@@
We compare PRISM and baselines in two settings: personalized T2I generation and direct image inversion, and we will elaborate on the task definitions in their corresponding sections below. For personalized T2I generation, we use a maximum budget of $40$ and report the quantitative results with $N=10,K=4$. For direct image inversion, we use a maximum budget of $30$ and report the quantitative results with $N=6,K=5$.
Baselines
@@
Evaluation Metrics
We evaluate the prompt interpretability using mean negative log-likelihood (NLL) calculated from Mistral 7B (Jiang et al., 2023). For image quality evaluation, we mainly measure the CLIP image similarity score (CLIP-I) to quantify the difference between the generated images and the reference images. Following Ruiz et al. (2023), we also use DINO V2 (Oquab et al., 2024) embedding similarity to calculate the object-sensitive image similarity for the personalized T2I generation task. We chose CLIP-ViT-L-14 and DINO-V2-Base as the base models. For Dall-E 2 and Dall-E 3, we also compare the number of times each method fails to pass its black-box safeguard. More failures indicate a higher potential to produce unsafe prompts. For each prompt, we allow $5$ attempts before counting it as a failure.
@@ 4.2 Personalized Text-to-Image Generation
We first demonstrate PRISM’s ability to find human-interpretable and accurate prompts to describe certain objects or styles in the task of personalized T2I generation. Given a set of reference images that depict the same concept (such as objects and style), this task requires the T2I model to synthesize images in new contexts while maintaining the original concept.
@@
Datasets
We use DreamBooth dataset (Ruiz et al., 2023) to quantitatively compare the performance in personalized T2I generation. DreamBooth dataset contains 30 daily objects and each subject has 4-6 images. For each subject, we adopt the 25 prompt templates curated by DreamBooth to create varying contexts and scenarios to test the fidelity of the subject representation in diverse settings. We generate 4 images for each subject and template combination with open sourced T2I models, and 1 image for each combination with closed sourced T2I models. For Textual Inversion, we follow the original setting to fill the templates and for all the other methods, we use the class noun to fill the template and the output prompts that describe these concepts serve as suffixes.
We also qualitatively demonstrate the ability to represent a certain artistic style using Wikiart dataset (Tan et al., 2019). We use three images from each artist as reference images. To create diverse scenes, we follow He et al. (2024) and use descriptive prompts from PartiPrompts (Yu et al., 2022) as prefixes to the output prompts similar to the previous setting.
DreamBooth Dataset Results
Table 1 and Figure 3 respectively show the quantitative and qualitative results on the DreamBooth dataset. As we can observe, PRISM achieves the best performance across the board except for the image similarity metrics for SDXL-Turbo.
In terms of object fidelity, we find PRISM to constantly achieve accurate depiction of the target subject while the baselines sometimes struggle to capture all fine-grained details like the colors of the animals and the shape of the shoe sole. And out of the four training-free methods we experiment with, PRISM is the only one that can attempt to tackle complicated objects such as the red monster toy and the dog-shaped backpack when all the other methods fail to generate even remotely similar objects. Due to the nature of their methodologies, BLIP-2 and CLIP-Interrogator also capture the background and other irrelevant elements in the scene when describing the objects. However, unlike our method, where we can directly specify the tasks and the judging criteria in the system prompts of the VLMs, there is no simple way to automatically filter out those irrelevant elements in BLIP-2 and CLIP-Interrogator’s outputs. Even though Textual Inversion obtains marginally higher CLIP-I scores and DINO scores with SDXL-Turbo, it requires a lot more modeling assumptions than our method, and the new embeddings it learns are not transferable – not even to SD 2.1.
PRISM is the only method in our experiments that can produce fully human-readable prompts while providing enough relevant details. In particular, we can observe that PEZ renders completely indecipherable texts, BLIP-2 only describes the general scene but fails to mention any visual details and textual inversion is entirely not interpretable since it produces soft embeddings. Since CLIP-Interrogator combines the results from BLIP-2 and a CLIP search, it improves the interpretability over PEZ-like gradient search-only method. However, it still falls short in terms of human readability in comparison to our method.
When transferring the output prompts to black-box T2I models, our method shows even larger advantages over the baselines. We also observe that our method produces the fewest unsafe prompts judged by Dall-E safeguards, while the baselines fail to pass the safeguard up to $16.7\%$ of the time.
Wikiart Results
In Figure 4, we show a qualitative comparison between our method and baselines on the Wikiart dataset. We find that our method is capable of precisely identifying the genres, eras, and sometimes even the names of the artists when describing the style of the reference artworks. On the other hand, the baselines fail to recognize these crucial keywords, even when they have access to a pre-collected bank of words that is supposed to provide accurate descriptions of art styles. In addition, PRISM can provide other fine-grained details such as pen strokes style and color palettes in a human-interpretable way to better assist the generation of the target style.
@@ 4.3 Direct Image Inversion
To demonstrate the versatility of our method, we also compare PRISM the baselines in the task of direct image inversion. In this task, the goal is to directly find the prompt that can exactly generate the input image. Here the number of reference images is $M=1$ and we aim to capture all aspects of the image, including the subjects, background, theme, style, and other details in the scene.
Datasets
We use images from the DiffusionDB dataset (Wang et al., 2022) for the direct image inversion task. This dataset includes a wide variety of image pairs generated by Stable Diffusion and we choose a random sample of 100 images from the large_random_10k split on Huggingface.
Results
As shown in Table 2, we immediately see a significant improvement in the human-interpretability of inverted prompts using PRISM. While expected for methods, such as PEZ, which has no language prior, we also find that our method finds text that more closely aligns with a learned distribution of English language text (i.e. lower NLL) than CLIP-Interrogator and BLIP2.
When comparing the image quality, we first note that because all images in DiffusionDB are generated by Stable Diffusion, which is exactly the model design space of CLIP-Interrogator and PEZ, it gives significant modeling assumption advantages to these baselines over our method when testing on Stable Diffusion models. This advantage enables relatively high performance for these baselines on Stable Diffusion models, but it does not transfer well into other closed-sourced models. In fact, we can even observe that CLIP-Interrogator generates the highest quality images with SD 2.1, which is the weakest model in this comparison, and the lowest quality images with Dall-E 3, which is the strongest T2I model in this table. This phenomenon indicates that the design choices of CLIP-Interrogate and PEZ are heavily overfitted on Stable Diffusion, and have poor generalizability to other models. On the other hand, the prompts produced by PRISM generalize significantly better than the baselines and we achieve better results with more powerful T2I models. When compared with the other VLM baseline Liu et al. (2024), with which a thorough comparison is included in Section F, PRISM performs it in almost all metrics, indicating a superior algorithmic design.
@@ 4.4 Ablation Study
Table 3: Comparison with GPT-4V in both personalized T2I generation and direct image inversion experiments.
@@
Comparison with GPT-4V
While we can use any VLM (which we demonstrate in Appendix C.2), it is nonetheless useful to understand what benefits PRISM  adds to an already capable foundation model like GPT-4V. Therefore, we compare our method with GPT-4V’s zero-shot performance with the same system prompts for both tasks on SDXL-Turbo. We can see in Table 3 that PRISM  consistently outperforms GPT-4V’s zero-shot performance, although the latter is already compelling. In Figure 6, we can also observe that qualitatively GPT-4V can capture the high-level semantics of the reference images but still misses fine-grained details.
Figure 7: Ablation study on different numbers of streams $N$ with fixed $K=5$.
Figure 8: Ablation study on different numbers of iterations $K$ with fixed $N=3$.
Effect of Budgets
Next we take a closer look at the effect of increasing the budget in PRISM. Figure 8 and 8 show the effect of increasing the number of streams $N$ and the number of iterations $K$ respectively. We observe that when increasing $N$ and keeping $K$ fixed, we can obtain steady performance improvements in both human readability and prompt accuracy. When increasing $K$ and keeping $N$ fixed, although we do not observe a monotonic relationship between the performance and $K$, we can still notice a general upward trend in prompt accuracy. In Appendix D, we discuss the trade-off between $N$ and $K$ to better inform practitioners how to choose these hyperparameters. Besides adjusting the budget, one can also use cheaper or open-sourced VLMs to lower the cost. In Appendix C.2, we experiment with GPT-4o-mini and IDEFICS2 (Laurençon et al., 2024), two significantly cheaper and smaller VLMs to demonstrate the cost-flexibility of PRISM.
4.5 Prompt Editing
Figure 9: Prompt editing demonstration with Midjourney.
Because the prompts produced by PRISM is very human-interpretable, after obtaining a prompt from the reference images, one can easily modify the output prompts to change attributes in their desired generated images. Figure 9 demonstrates an examples of prompt editing with PRISM on Midjourney. With simple and intuitive prompt edits, we are able to change specific attributes of the images while keeping the other components in the scene relatively unchanged.
5 Conclusion
In this paper, we propose PRISM, an algorithm that automatically creates human-interpretable and accurate text prompts for text-to-image generative models, based on visual concepts provided by reference images. Our method iteratively refines the sampling distribution of the text prompt via VLM in-context learning and is capable of creating prompts that are transferable to any T2I models, including black-box platforms like Dall-E and Midjourney. We hope our work also encourages researchers, particularly those in non-LLM fields, to consider how the advancements in LLMs can offer simple yet effective solutions to problems that pre-LLM methods have struggled to address.
Ethic Statement
Just as LLMs are suceptible to being jailbroken or adversarially manipulated by malicious actors (Zou et al., 2023), our method may also be vulnerable to malicious intent, potential bias, or limitations in the base models. Therefore, we will implement necessary safeguards upon the public release of our code and are committed to keep up with future advancements in improving the safety of our method.
Acknowledgements
This work was supported in part of the ONR grant N000142312368. We also thank Sony AI for sponsoring part of this work. A.R., H.H., and G.P. are supported by the NSF Institute for CORE Emerging Methods in Data Science (EnCORE). A.R. is also supported by an ASSET Amazon AWS Trustworthy AI Fellowship. Y.J. is supported by Google PhD Fellowship. We thank Murtaza Dalal, Martin Ma, Lawrence Jiang, Aakash Lahoti, Aviv Bick, Issac Liao, Sukjun Hwang, Kevin Li and Ricardo Buitrago for their feedback on this paper.
Appendix A Additional Experiment Details
In this section, we provide further details about the implementation of our experiments. For all quantitative analysis that uses Stable Diffusion based model, we generate four images for each combination of prefixes and prompts. For all experiments with Dall-E based model, we generate one image per combination. In the DreamBooth dataset experiment, we also replace the class noun for “stuffed animal” with “toy” to obtain fair comparisons with textual inversion, which can only take a single token as the initialization token. We use OpenCLIP-ViT-H-14 trained on LAION2B (Schuhmann et al., 2022) for both CLIP-Int and PEZ and use Blip2-Flan-T5-XL for both CLIP-Int and BLIP-2.
To simplify the implementation, we only keep a chat history length of $3$ and use the length of the prompt as an approximation of the log-likelihood for the final prompt selection. When evaluating the judge scores $\mathbf{D}(x,\hat{x})$ in PRISM iterations, we shuffle the reference images when $M>1$. The judge score is rescaled into a range from 0 to 10. For direct image inversion, we re-evaluate the top $5$ candidates twice and tally the scores with in-iteration scores to make the final decision. For personalized T2I generation, we re-evaluate once for each reference image and use the average score to select the output.
We provide the demo code of our method here.
Appendix B Designing System Prompts
System prompting is the standard way to condition a general purpose LLM for specific tasks of request. The key idea is that, before the conversation starts, the LLM receives a tailored message, the system prompt, that provides the contexts, conversation scenario settings, formats and other guidelines as the prefix of the entire conversation ahead. In this section, we elaborate on the design of the system prompts for the prompt engineer assistant ${\mathbf{F}}$ and the judge ${\mathbf{D}}$. We also provide the full system prompts used in all of our experiments at the end of this paper in Section G and in our demo code.
B.1 Prompt Engineer Assistant ${\mathbf{F}}$
To design the system prompts for the prompt engineer assistant ${\mathbf{F}}$, we follow Chao et al. (2023) and include the following components in the system prompt of ${\mathbf{F}}$.
Setting
We first set up the scenarios and assign a role for the LLM to perform better on the specific task of choice. The setting paragraphs start with “You are a helpful prompt engineer assistant. You are free to generate sentences that do NOT follow English grammar. You must obey all of the following instructions.” and continue with the specific description of the task and the objective. We also inform the assistant that it is expected to iterate and refine the prompts it generates throughout the conversation.
Format
We then provide the guidelines for formatting the inputs and the outputs of the assistant. We describe what are expected in the inputs at each iteration and the content required in the outputs. We also provide descriptions of the meanings of each input and output components. More specifically, we inform the assistant that the inputs consist of three parts: a generated image, a reference images, and a visual similarity score, and that the assistant is expected to generate both the improvement to refine the previous prompt and the next new prompt. All generated text is formatted in JSON.
Examples
B.2 Judge ${\mathbf{D}}$
Appendix C Additional Results
C.1 Additional Qualitative Results
In Figure 10, 11, 12, 13 and 14, we provide additional qualitative showcases for subject-driven personalized T2I generation, style-driven personalized T2I generation, direct image inversion and prompt editing. We also provide an example of the iteration and refinement process as a conversation between all three components in PRISM in Figure 15.
Figure 13: Qualitative examples of the direct image inversion task.
C.2 Flexible Model Choices
To further demonstrate the effectiveness and flexibility of PRISM, we conduct further experiments with diverse model choices for prompt engineer assistant ${\mathbf{F}}$, T2I generator ${\mathbf{G}}$ and judge ${\mathbf{D}}$. In Table 4, we first provide additional validation of the flexibility for choosing different VLM as base models for ${\mathbf{F}}$ and ${\mathbf{D}}$ using (1) a significantly smaller open-source model IDEFICS2 (Laurençon et al., 2024), with IDEFICS2-8b-chatty as ${\mathbf{F}}$ and IDEFICS2-8b as ${\mathbf{D}}$ and (2) a significantly smaller and cheaper closed-source model GPT-4o-mini as both ${\mathbf{F}}$ and ${\mathbf{D}}$. While there is some expected performance drop compared to GPT-4V, PRISM still delivers very competitive results and notably maintains human-readability and generalizability, particularly with the closed-sourced model Dall-E 2. This aligns with our previous conclusion and underscores PRISM’s adaptability across various computational environments.
Table 4: Additional experimental results with different VLMs as ${\mathbf{F}}$ on DreamBooth dataset.
[…]
TI (SDXL) - 0.771 0.504 - - -
CLIP-Int 4.361 0.756 0.490 0.711 0.464 13.3%
BLIP2 4.378 0.729 0.456 0.707 0.430 6.9%
PEZ 6.188 0.722 0.418 0.676 0.389 16.7%
PRISM (IDEFICS2) 3.047 0.739 0.468 0.721 0.453 6.7%
PRISM (GPT-4o-mini) 3.498 0.768 0.493 0.730 0.475 6.7%
PRISM (GPT-4V) 3.466 0.770 0.499 0.734 0.482 6.9%
We also experiment a different T2I Generator ${\mathbf{G}}$ to showcase the transferability of the prompts generated by PRISM. Figure 16 shows qualitative examples of PRISM prompts with Dall-E 2 as the Generator ${\mathbf{G}}$ for personalized T2I generation and the images generated from those prompts using SDXL-Turbo, Dall-E 3 and Midjourney. Our method is capable of producing human-interpretable and accurate prompts for both subject-driven T2I personalization and style-driven T2I personalization with this new Generator ${\mathbf{G}}$.
C.3 Additional Applications
In this section, we demonstrate two additional applications of PRISM, prompt distillation and multi-concept generation.
PRISM is particularly well-suited for multi-concept generation due to the human readability of its generated prompts. This feature allows for easy identification and composition of different components within a scene, enabling intuitive control over multi-concept results. Unlike PEZ, which does not provide explicit control over which part of the prompt corresponds to specific aspects of the image (e.g., content or style), PRISM allows for much clearer and more direct manipulation, which we demonstrate in Figure 17
Figure 17: Qualitative demonstration of multi-concept generation with PRISM.
Unlike PEZ, which requires an additional optimization process to generate distilled prompts, PRISM leverages its highly interpretable prompts and the capabilities of LLMs to simplify prompts effectively. By using in-context learning and straightforward instructing a VLM (in this case a GPT-4o model) with the prompt “Here is a prompt to this image with a text-to-image model, make it more concise (less than $<$ token length constraint $>$ tokens) but keep all the descriptive details”, we achieve concise, distilled prompts without additional computational overhead, as we can observe in Figure 18.
Figure 18: Qualitative demonstration of prompt distillation with PRISM.
Appendix D Additional Ablation Study
In this section, we provide a more detailed ablation study on each component of the PRISM framework. In particular, we demonstrate the trade-off between the number of streams $N$ and the number of iterations $K$, compare a non-VLM judge (a CLIP judge) against our choice of a LLM judge (GPT-4V Judge), and also the effect of the existence of the Judge ${\mathbf{D}}$ and re-evaluation.